when 4 liters are added to a tank that is 1/3 full the tank is then 1/2 full. what is the capacity of the tank?
Answer by rfer(16322)

1/2-1/3=
3/6-2/6=1/6
-----------
4/(1/6)=
4*6/1=24 liters